# Beyond Consensus: Mitigating the Agreeableness Bias in LLM Judge Evaluations

### v1

@@ 1 Introduction
This has caused developers to turn to automated evaluation. While benchmarking is straightforward for tasks with unambiguous correctness criteria, such as multiple-choice questions in MMLU (Hendrycks et al., 2021) or coding problems with unit-test oracles in HumanEval (Chen et al., 2021), benchmarking is significantly more challenging for a new class of open-ended problems where multiple distinct outputs can be correct, such as generating feedback for incorrect student programs from the IntroPython dataset (Sahai et al., 2023). In this paper, we investigate the challenges of using LLMs as judges for this emerging problem class, which is active research area.
The state-of-the-art automated benchmarking approach, “LLM-as-a-judge” (Zheng et al., 2023), suffers from high variance across different validator models, making it difficult to select a reliable judge (Wang et al., 2023). More critically, as observed in prior work (Thakur et al., 2024), we found even the latest LLM models exhibit a strong positive bias, i.e. they are proficient in identifying correct answers, but they perform poorly in identifying incorrect ones. Our analysis reveals a low True Negative Rate (TNR) across all models (§3), which leads to an overestimation of model precision. This low TNR undermines the reliability of LLM-based evaluation, especially since the proportion of invalid outputs is typically small. Furthermore, popular benchmarking frameworks such as Elo scoring (Zheng et al., 2023; Chiang et al., 2024), while cleverly using human preferences for relative model comparisons, do not reflect absolute model performance or evaluator reliability. As we can see in Figure 1, even high Elo-ranked models can exhibit poor TNR, suggesting that Elo rankings alone are insufficient to decide on the best model for a task.
@@
Figure 1: Correlation between LLM Elo ratings (Chiang et al., 2024) and their performance as generators and validators on high-school programming feedback.
@@
2. New dataset and code for subjective task: To support reproducible research, we release an extended dataset building on Sahai et al. (2023), with 366 buggy programs, feedback generated by 14 LLMs, and validation judgments from all 14 models, along with human annotations for 6 of the generators [^1] Our code and data is available at https://github.com/ai-cet/llm-judge-calibration.
@@ 2 Problem Setting and Motivation
As a motivating example, consider a typical introductory programming course (CS1). Given a problem description, test cases, and a student’s buggy program, the task is to generate feedback that helps the student identify and fix their mistakes. In our work, $\mathcal{T}$ is a dataset of $366$ incorrect Python program submissions for $69$ different assignments (Sahai et al., 2023). Task $t$ is formally defined as: “Given a student’s incorrect code, problem description, and test cases, generate valid feedback to help the student fix their mistake(s).”
@@
Ground Truth Data. Our evaluation is based on the high school Python programming assignment benchmark (Sahai et al., 2023), which includes human annotations of feedback generated by GPT 4. We extend this by annotating the outputs of five additional LLMs: Google’s Gemini 1.5-Pro, Anthropic’s Claude Opus 3, OpenAI’s GPT 4o, Qwen Coder Plus, and Deepseek V2.5. For each of the $366$ tasks, we prompted each LLM to generate feedback. All LLM prompts are included in the supplementary materials. Since feedback is granular and often tied to specific lines of code, a single submission can elicit multiple feedback items.
@@ 3 LLM as Validators
LLMs are increasingly used to evaluate the output of other LLMs (Zheng et al., 2023). This raises a natural question: How reliable are LLMs at validating each other’s outputs?
@@ 4 Ensemble of LLM Validators
To mitigate the unreliability of individual LLMs, a natural approach is to employ an ensemble (Lu et al., 2024; Jiang et al., 2023). The standard strategy is to use a Simple Majority consensus, which assigns a label as “valid” if the majority of validators agree and “invalid” otherwise. With a threshold of $8$ (out of $14$) validators, this strategy reduces the maximum error on our six annotated generators from 17.617.6% for the worst individual LLM to 14.814.8% by using an ensemble. However, this approach is highly sensitive to missing data. After applying data repair techniques to reduce missing values from 9.79.7% to 3.53.5%, the maximum error for Simple Majority drops to 4.84.8%, highlighting its dependency on data completeness. These repairs involved three main steps: first, label standardization to unify validator outputs (e.g., mapping “incorrect” to ‘invalid’); second, fuzzy string matching (Levenshtein distance) to align feedback items; and third, for feedback associated with line number ranges (e.g., ‘2-3’), we match the starting line number if an exact range match is not found.
@@ 5.1 Methodology
We solve this optimization problem using the L-BFGS-B algorithm (Zhu et al., 1997), a quasi-Newton method that efficiently handles a large number of free variables by approximating the Hessian matrix using limited memory, with box constraints to ensure all estimated parameters remain within $[0,1]$.
@@ 6 Related Work
LLM-as-Judge. The usage of LLMs for evaluating other LLMs is now widespread (Gu et al., 2024). Prior work has identified various biases, such as preferences for position, verbosity, and a general leniency or “self-enhancement” bias where LLMs rate their own outputs favorably (Zheng et al., 2023). Our work differs by framing this leniency as a quantifiable “agreeableness bias”, characterized by an imbalance between high TPR and low TNR. While relatively ranking LLMs in terms of strength is popular (Chiang et al., 2024), we make use of a binary (True/False) setup instead as relative ranking can obscure the absolute correctness or precision of generator outputs, as seen in Figure 1(a).
Ensemble Methods. Ensemble approaches are a natural solution to mitigate individual model biases. While simple majority voting is common, recent work has explored more sophisticated strategies like generative fusion (Jiang et al., 2023) and selecting diverse sub-ensembles (Tekin et al., 2024). However, these methods remain fundamentally constrained by the systematic biases of their constituent models. As demonstrated in §4, standard voting is insufficient for addressing the low TNR that characterizes LLM judges, and even our optimized minority-veto strategy is limited by this inherent flaw.
Calibration and Latent Trait Models. Our regression framework is conceptually related to classic models for aggregating noisy labels, such as the EM Algorithm (Dawid & Skene, 1979) and Item Response Theory (IRT) (Lord, 2012). While IRT is powerful for relative ranking, its latent variables do not directly map to an absolute, interpretable metric like precision. Our model is tailored to this estimation task by explicitly modeling generator precision ($g_{i}$) as a variable, similar to calibration techniques, which aim to align model scores with correctness probabilities (Guo et al., 2017). However, existing methods typically address general overconfidence rather than the specific, systematic low-TNR problem that we identify. Our regression-based approach differs by using a ground-truth calibration set to explicitly model and correct for this bias, shifting from general confidence scaling to targeted bias correction. We refer the reader to our Appendix E for a more detailed discussion.
@@ Acknowledgements
This research is supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-009-AICET). This project was supported by Google.org with a Google Cloud Platform (GCP) credit award.
@@ B.3 Repair Missing Values
We repair some of the missing values in the dataset by applying fuzzy matching on the feedback, line-number and labels. As explained in our main paper, a large number of LLM validator outputs could not be aligned with the LLM generated feedback due to issues in reproducing the original feedback, line number, or binary label (valid/invalid). Instead of an exact match, we make use the FuzzyWuzzy library (Cohen, 2011) to match the feedback referenced by the LLM validator with LLM generator’s feedback. If the match score is above 85% threshold, we declare the match a success. Similarly, we fix issues with line numbers, where the validator could have used an incorrect key (e.g, “line_num” instead of ‘line_number‘) or a different format (e.g., “line 3” instead of “3”). We also fix the binary labels, where the validator could have invented a new label (e.g., ‘’‘partially valid” instead of the expected “valid” and “invalid”). After applying these repairs, the missing values reduce drastically from 9.7% to 3.5%. Following which, Simple Majority consensus ($m=8$ valid voting ensemble) had a major improvement in its max error rate from 14.8% to 4.8%.
@@ Appendix D Regression Additional Results
We solve the optimization problem using the BFGS algorithm (Fletcher, 2000) with a convergence tolerance of $10^{-6}$ on the prediction loss. The parameters are initialized as follows: generator precisions $\widehat{g_{i}}$ are set to the mean of their corresponding row in the prediction matrix $\mathrm{P}$, while validator TPRs $\widehat{v}_{j}^{+}$ and TNRs $\widehat{v}_{j}^{-}$ are initialized to 1.0. To mitigate the risk of local minima, we perform 10 runs with small random offsets added to the initial values and select the run with the lowest training error.
@@ Appendix E Related Work - Expanded
Human Evaluation. Human evaluation has traditionally been the gold standard for assessing language model performance, especially for inherently subjective tasks such as translation, summarization, and open-ended question answering (Brown et al., 2020). In these evaluations, human judges rate the quality of model outputs based on specific criteria, providing reliable insights into the model’s capabilities. Another approach involves collecting human preferences between two or more model outputs in head-to-head comparisons (Zheng et al., 2023). However, these approaches lack scalability, particularly with the rapid release of newer models requiring frequent evaluation.
@@
LLM-as-a-Judge. The use of Large Language Models as evaluators has gained significant traction for assessing open-ended text generation quality, often achieving over $80\%$ agreement with human judgments (Zheng et al., 2023). However, recent work has revealed critical vulnerabilities within these systems. Studies indicate that even state-of-the-art LLMs suffer from poor recall, failing to identify the majority of errors made by other LLMs on subjective tasks (Kamoi et al., 2024). LLM judges has earlier been reported to exhibit systematic biases including positional bias where response order and persuasive repetitive language influences judgments (Wang et al., 2023; Pezeshkpour & Hruschka, 2024); and more critically, what we term “agreeableness bias,” or a tendency to accept invalid outputs as correct. Simple adversarial inputs that suggest deep reasoning such as “Thought process:” or even non-word symbols can trigger false positive rewards exceeding $80\%$ for state-of-the-art models including GPT 4o and Claude Sonnet 4 (Zhao et al., 2025). We are the first to show with data that the TNR for existing LLMs are often below $25\%$.
@@
An alternative to using off-the-shelf LLMs as judges is to fine-tune a verifier model, which has been shown to improve reliability on datasets like GSM8K (Cobbe et al., 2021). However, this approach is costly, prone to overfitting, and yields limited benefits with scarce training data. Furthermore, a verifier trained on one dataset often lacks generalizability, requiring retraining for new evaluation tasks.
@@
Ensemble Methods. Ensemble approaches represent a natural solution to mitigate individual LLM biases by leveraging diverse model strengths. While simple jury voting schemes are common, recent work has explored more sophisticated strategies, as reviewed in Chen et al. (2025). LLM-BLENDER (Jiang et al., 2023) combines pairwise ranking with generative fusion, using specialized comparison methods to distinguish candidate outputs and generate a unified high quality response. The Mixture-of-Reasoning-Experts (MORE) framework (Si et al., 2023) creates specialized experts for different types of reasoning and leverages inter-expert agreement for both answer selection and abstention, achieving superior generalizability and question answering performance.
@@
Mirror-Consistency Huang et al. (2024a) enhances Self-Consistency by incorporating a “reflective mirror” that analyzes inconsistencies among multiple generations rather than simply aggregating them. This approach shows improvements in both reasoning accuracy and confidence calibration by enabling models to learn from discrepancies in their reasoning pathways. Recent advances include Speculative Ensemble methods that accelerate LLM by generating and verifying tokens in parallel without sacrificing performance (Fu et al., 2025), and diversity-optimized ensemble techniques that use focal diversity metrics to select optimal sub-ensembles from larger model pools (Tekin et al., 2024).
@@
Calibration and Bias Correction. Calibration techniques aim to ensure model confidence scores accurately reflect correctness probability. Traditional approaches like temperature scaling have been extended through adaptive means, but these primarily address general overconfidence rather than systematic evaluation biases (Guo et al., 2017). In the LLM domain, selective generation based calibration methods leverage consistency across multiple generations to identify unreliable results (Lin et al., 2023).
However, existing calibration methods focus on confidence scaling rather than addressing the fundamental issue we identify: systematic tendencies of LLM judges to agree with invalid outputs. Recent work on post-hoc reward calibration attempts to estimate bias terms using local reward averaging approaches (Huang et al., 2024b), but these methods do not specifically target the low TNR problem that compromises evaluation reliability. Our regression-based approach differs by explicitly modeling generator-validator interactions and using ground-truth data to calibrate for specific biases, representing a shift from general confidence calibration to targeted bias correction for evaluation systems.

### v2

@@ 1 Introduction
This has caused developers to turn to automated evaluation. While benchmarking is straightforward for tasks with unambiguous correctness criteria, such as multiple-choice questions in MMLU (mmlu2021) or coding problems with unit-test oracles in HumanEval (humanEval), benchmarking is significantly more challenging for a new class of open-ended problems where multiple distinct outputs can be correct, such as generating feedback for incorrect student programs from the IntroPython dataset (sahai2023improving). In this paper, we investigate the challenges of using LLMs as judges for this emerging problem class, which is active research area.
The state-of-the-art automated benchmarking approach, “LLM-as-a-judge” (zheng2023LLMJudge), suffers from high variance across different validator models, making it difficult to select a reliable judge (wang2023large). More critically, as observed in prior work (thakur2024judging), we found even the latest LLM models exhibit a strong positive bias, i.e. they are proficient in identifying correct answers, but they perform poorly in identifying incorrect ones. Our analysis reveals a low True Negative Rate (TNR) across all models (§3), which leads to an overestimation of model precision. This low TNR undermines the reliability of LLM-based evaluation, especially since the proportion of invalid outputs is typically small. Furthermore, popular benchmarking frameworks such as Elo scoring (zheng2023LLMJudge; chiang2024chatbot), while cleverly using human preferences for relative model comparisons, do not reflect absolute model performance or evaluator reliability. As we can see in Figure 1, even high Elo-ranked models can exhibit poor TNR, suggesting that Elo rankings alone are insufficient to decide on the best model for a task.
@@
Figure 1: Correlation between LLM Elo ratings (chiang2024chatbot) and their performance as generators and validators on high-school programming feedback.
@@
2. New dataset and code for subjective task: To support reproducible research, we release an extended dataset building on sahai2023improving, with 366 buggy programs, feedback generated by 14 LLMs, and validation judgments from all 14 models, along with human annotations for 6 of the generators [^1] Our code and data is available at https://github.com/ai-cet/llm-judge-calibration.
@@ 2 Problem Setting and Motivation
As a motivating example, consider a typical introductory programming course (CS1). Given a problem description, test cases, and a student’s buggy program, the task is to generate feedback that helps the student identify and fix their mistakes. In our work, $\mathcal{T}$ is a dataset of $366$ incorrect Python program submissions for $69$ different assignments (sahai2023improving). Task $t$ is formally defined as: “Given a student’s incorrect code, problem description, and test cases, generate valid feedback to help the student fix their mistake(s).”
@@
Ground Truth Data. Our evaluation is based on the high school Python programming assignment benchmark (sahai2023improving), which includes human annotations of feedback generated by GPT 4. We extend this by annotating the outputs of five additional LLMs: Google’s Gemini 1.5-Pro, Anthropic’s Claude Opus 3, OpenAI’s GPT 4o, Qwen Coder Plus, and Deepseek V2.5. For each of the $366$ tasks, we prompted each LLM to generate feedback. All LLM prompts are included in the supplementary materials. Since feedback is granular and often tied to specific lines of code, a single submission can elicit multiple feedback items.
@@ 3 LLM as Validators
LLMs are increasingly used to evaluate the output of other LLMs (zheng2023LLMJudge). This raises a natural question: How reliable are LLMs at validating each other’s outputs?
@@ 4 Ensemble of LLM Validators
To mitigate the unreliability of individual LLMs, a natural approach is to employ an ensemble (lu2024merge; jiang2023llm). The standard strategy is to use a Simple Majority consensus, which assigns a label as “valid” if the majority of validators agree and “invalid” otherwise. With a threshold of $8$ (out of $14$) validators, this strategy reduces the maximum error on our six annotated generators from 17.617.6% for the worst individual LLM to 14.814.8% by using an ensemble. However, this approach is highly sensitive to missing data. After applying data repair techniques to reduce missing values from 9.79.7% to 3.53.5%, the maximum error for Simple Majority drops to 4.84.8%, highlighting its dependency on data completeness. These repairs involved three main steps: first, label standardization to unify validator outputs (e.g., mapping “incorrect” to ‘invalid’); second, fuzzy string matching (Levenshtein distance) to align feedback items; and third, for feedback associated with line number ranges (e.g., ‘2-3’), we match the starting line number if an exact range match is not found.
@@ 5.1 Methodology
We solve this optimization problem using the L-BFGS-B algorithm (lbfgsb), a quasi-Newton method that efficiently handles a large number of free variables by approximating the Hessian matrix using limited memory, with box constraints to ensure all estimated parameters remain within $[0,1]$.
@@ 6 Related Work
LLM-as-Judge. The usage of LLMs for evaluating other LLMs is now widespread (gu2024survey). Prior work has identified various biases, such as preferences for position, verbosity, and a general leniency or “self-enhancement” bias where LLMs rate their own outputs favorably (zheng2023LLMJudge). Our work differs by framing this leniency as a quantifiable “agreeableness bias”, characterized by an imbalance between high TPR and low TNR. While relatively ranking LLMs in terms of strength is popular (chiang2024chatbot), we make use of a binary (True/False) setup instead as relative ranking can obscure the absolute correctness or precision of generator outputs, as seen in Figure 1(a).
Ensemble Methods. Ensemble approaches are a natural solution to mitigate individual model biases. While simple majority voting is common, recent work has explored more sophisticated strategies like generative fusion (jiang2023llm) and selecting diverse sub-ensembles (tekin2024llm). However, these methods remain fundamentally constrained by the systematic biases of their constituent models. As demonstrated in §4, standard voting is insufficient for addressing the low TNR that characterizes LLM judges, and even our optimized minority-veto strategy is limited by this inherent flaw.
Calibration and Latent Trait Models. Our regression framework is conceptually related to classic models for aggregating noisy labels, such as the EM Algorithm (dawid1979maximum) and Item Response Theory (IRT) (lord2012applications). While IRT is powerful for relative ranking, its latent variables do not directly map to an absolute, interpretable metric like precision. Our model is tailored to this estimation task by explicitly modeling generator precision ($g_{i}$) as a variable, similar to calibration techniques, which aim to align model scores with correctness probabilities (guo2017calibration). However, existing methods typically address general overconfidence rather than the specific, systematic low-TNR problem that we identify. Our regression-based approach differs by using a ground-truth calibration set to explicitly model and correct for this bias, shifting from general confidence scaling to targeted bias correction. We refer the reader to our Appendix E for a more detailed discussion.
@@ Acknowledgements
This research is supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-009-AICET). This project was supported by Google.org with a Google Cloud Platform (GCP) credit award. We would like to thank Xu Yin and his team at the NUS Research Institute for their valuable discussions and support in annotating outputs from DeepSeek and Qwen Coder, which contributed to this work.
@@ B.3 Repair Missing Values
We repair some of the missing values in the dataset by applying fuzzy matching on the feedback, line-number and labels. As explained in our main paper, a large number of LLM validator outputs could not be aligned with the LLM generated feedback due to issues in reproducing the original feedback, line number, or binary label (valid/invalid). Instead of an exact match, we make use the FuzzyWuzzy library (fuzzywuzzy) to match the feedback referenced by the LLM validator with LLM generator’s feedback. If the match score is above 85% threshold, we declare the match a success. Similarly, we fix issues with line numbers, where the validator could have used an incorrect key (e.g, “line_num” instead of ‘line_number‘) or a different format (e.g., “line 3” instead of “3”). We also fix the binary labels, where the validator could have invented a new label (e.g., ‘’‘partially valid” instead of the expected “valid” and “invalid”). After applying these repairs, the missing values reduce drastically from 9.7% to 3.5%. Following which, Simple Majority consensus ($m=8$ valid voting ensemble) had a major improvement in its max error rate from 14.8% to 4.8%.
@@ Appendix D Regression Additional Results
We solve the optimization problem using the BFGS algorithm (fletcher2000practical) with a convergence tolerance of $10^{-6}$ on the prediction loss. The parameters are initialized as follows: generator precisions $\widehat{g_{i}}$ are set to the mean of their corresponding row in the prediction matrix $\mathrm{P}$, while validator TPRs $\widehat{v}_{j}^{+}$ and TNRs $\widehat{v}_{j}^{-}$ are initialized to 1.0. To mitigate the risk of local minima, we perform 10 runs with small random offsets added to the initial values and select the run with the lowest training error.
@@ Appendix E Related Work - Expanded
Human Evaluation. Human evaluation has traditionally been the gold standard for assessing language model performance, especially for inherently subjective tasks such as translation, summarization, and open-ended question answering (brown2020language). In these evaluations, human judges rate the quality of model outputs based on specific criteria, providing reliable insights into the model’s capabilities. Another approach involves collecting human preferences between two or more model outputs in head-to-head comparisons (zheng2023LLMJudge). However, these approaches lack scalability, particularly with the rapid release of newer models requiring frequent evaluation.
@@
LLM-as-a-Judge. The use of Large Language Models as evaluators has gained significant traction for assessing open-ended text generation quality, often achieving over $80\%$ agreement with human judgments (zheng2023LLMJudge). However, recent work has revealed critical vulnerabilities within these systems. Studies indicate that even state-of-the-art LLMs suffer from poor recall, failing to identify the majority of errors made by other LLMs on subjective tasks (kamoi2024evaluating). LLM judges has earlier been reported to exhibit systematic biases including positional bias where response order and persuasive repetitive language influences judgments (wang2023large; pezeshkpour2024large); and more critically, what we term “agreeableness bias,” or a tendency to accept invalid outputs as correct. Simple adversarial inputs that suggest deep reasoning such as “Thought process:” or even non-word symbols can trigger false positive rewards exceeding $80\%$ for state-of-the-art models including GPT 4o and Claude Sonnet 4 (zhao2025one). We are the first to show with data that the TNR for existing LLMs are often below $25\%$.
@@
An alternative to using off-the-shelf LLMs as judges is to fine-tune a verifier model, which has been shown to improve reliability on datasets like GSM8K (cobbe2021training). However, this approach is costly, prone to overfitting, and yields limited benefits with scarce training data. Furthermore, a verifier trained on one dataset often lacks generalizability, requiring retraining for new evaluation tasks.
@@
Ensemble Methods. Ensemble approaches represent a natural solution to mitigate individual LLM biases by leveraging diverse model strengths. While simple jury voting schemes are common, recent work has explored more sophisticated strategies, as reviewed in chen2025harnessing. LLM-BLENDER (jiang2023llm) combines pairwise ranking with generative fusion, using specialized comparison methods to distinguish candidate outputs and generate a unified high quality response. The Mixture-of-Reasoning-Experts (MORE) framework (si2023getting) creates specialized experts for different types of reasoning and leverages inter-expert agreement for both answer selection and abstention, achieving superior generalizability and question answering performance.
@@
Mirror-Consistency huang2024mirror enhances Self-Consistency by incorporating a “reflective mirror” that analyzes inconsistencies among multiple generations rather than simply aggregating them. This approach shows improvements in both reasoning accuracy and confidence calibration by enabling models to learn from discrepancies in their reasoning pathways. Recent advances include Speculative Ensemble methods that accelerate LLM by generating and verifying tokens in parallel without sacrificing performance (fu2025speculative), and diversity-optimized ensemble techniques that use focal diversity metrics to select optimal sub-ensembles from larger model pools (tekin2024llm).
@@
Calibration and Bias Correction. Calibration techniques aim to ensure model confidence scores accurately reflect correctness probability. Traditional approaches like temperature scaling have been extended through adaptive means, but these primarily address general overconfidence rather than systematic evaluation biases (guo2017calibration). In the LLM domain, selective generation based calibration methods leverage consistency across multiple generations to identify unreliable results (lin2023generating).
However, existing calibration methods focus on confidence scaling rather than addressing the fundamental issue we identify: systematic tendencies of LLM judges to agree with invalid outputs. Recent work on post-hoc reward calibration attempts to estimate bias terms using local reward averaging approaches (huang2024post), but these methods do not specifically target the low TNR problem that compromises evaluation reliability. Our regression-based approach differs by explicitly modeling generator-validator interactions and using ground-truth data to calibrate for specific biases, representing a shift from general confidence calibration to targeted bias correction for evaluation systems.